Find the potential at 25oC of the half cell:
Pt(s)/Cu2+(aq), (0.50 M), Cu+(aq), (0.052 M)
Eo = 0.153 V.
Select one:
a. 0.153 V
b. 0.095 V
c. 0.182 V
d. 0.211 V
e. 0.269 V
1
Expert's answer
2017-08-29T06:17:57-0400
E = E(0) + 0.059lg([Ox]/[Red])
E = 0.153 + 0.059lg(0.5/0.052) = 0.153 + 0.059*0.98 = 0.211 (V)